Benson Manufacturing is considering ordering electronic components from three different suppliers. The suppliers may differ in terms of quality in that the proportion or percentage of defective components may differ among the suppliers. To evaluate the proportion defective components for the suppliers, Benson has requested a sample shipment of 500 components from each supplier. The number of defective components and the number of good components found in each shipment is as follows. Supplier Component 10 35 Good 490 465 455 a. Formulate the hypotheses that can be used to test for equal proportions of defective components provided by the three suppliers. 1. Ho: P p2 p3 Choose correct answer from above choice HaNot all population proportions are equal b. Using a.05 level of significance, conduct the hypothesis test. What is the p-value? Use Table 12.4. The p-value is 1 less than or equal .05 τ What is your conclusion? Conclude that the three suppliers do not provide equal proportions of defective c c. Conduct a multiple comparison test to determine if there is an overall best supplier or if one supplier can be eliminated because of poor quality. Round ρ i decimal places. and difference to two decimal places. Round critical value to four Comparison ρί Difference Critical Value Significant Diff CV A vs B Avs C B vs C Yes No
Answer #1

Proportion of defectives components:

10 0.02 PA 10490

35 = 0.07 35 465

45 pc- + 455 0.09

Degree of freedom = (3-1)*(2-1) = 2

At alpha = 0.05 and df =2, critical value of chi square, chi^2_c = 5.991

Critical value for A vs B =

n B RA

/0.02( 1-0.02) . О.07(1-007) v .991 * V = 00019

Critical value for A vs C =

nc

/0,02(1-0,02) . О.09(1-0.09) v/5soi * V = 0.0349

Critical value for B vs C =

nc n B

0.07(1-0.07) 0.09(1-009) = V5.991 * = 0.0420 .

Comparison p_i p_j Difference n_i n_j Critical value Significant Diff> CV
A vs B 0.02 0.07 0.05 500 500 0.0319 Yes
A vs C 0.02 0.09 0.07 500 500 0.0349 Yes
B vs C 0.07 0.09 0.02 500 500 0.0420 No